# Structure theorems for braided Hopf algebras

### v1

@@ Abstract
The cross terms cancel, showing that $[a,b]=ab-ba$ is primitive.
@@
Consequently, the same argument that $P(A)$ is a Lie algebra fails in the braided context. A great deal of the technical work in this paper is developing the notion of a braided primitive algebra [^1] The term “braided Lie algebra” is already in use in multiple (inequivalent) forms; see section 10 for a review of some of these terms and their relation to the constructions in this paper., which encapsulates the algebraic properties of $P(A)$. This structure is encoded by the data of an action of a braided operad which we notate as $\operatorname{BrPrim}$ and define in Section 5.2.
@@
The primitives $P(A)$ sit inside an larger space $P_{\mathfrak{W}_{*}}(A)\subseteq A$ whose definition (in section 7.1) requires some care, but can loosely be described as “all of the elements of $A$ that become primitive after iteratively quotienting by decomposable primitives in $A$” (see Proposition 9.2 for a more precise description). It supports an algebraic structure we are calling [^2] This term is also used in the setting of $C^{*}$-algebras; these are different notions, and it should be clear from context which is meant. a Woronowicz algebra. The corresponding operad is denoted $\mathfrak{W}_{*}$; it is defined in section 5.5 in terms of the “quantum symmetrizer” which gives the relations in Nichols algebras. It is larger and even less well-understood than $\operatorname{BrPrim}$.
@@
If $L$ is a Woronowicz algebra equipped with a diagonal, there is a filtration on $U_{\mathfrak{W}_{*}}(L)$ by powers of the primitives $P(L)=P(U_{\mathfrak{W}_{*}}(L))$. Unlike the classical setting, it is possible that more primitives are produced in the associated graded Hopf algebra $U_{\mathfrak{W}_{*}}(L)^{\operatorname{gr}}$. We may, however, iterated this procedure, producing a graded Hopf algebra $U_{\mathfrak{W}_{*}}(L)^{(\infty)}$. Our analogue of the Poincaré-Birkhoff-Witt theorem is:
@@
We actually prove versions of Theorems 1 and 2 for a large class of operads $\mathcal{C}$; see Theorem 7.10 and Proposition 8.14 for precise statements. The Woronowicz ideal $\mathfrak{W}$ is the smallest operad $\mathcal{C}$ that we have identified for which these results hold, but still very large. It is very plausible that there is a class of simpler (or at least smaller) operads $\mathcal{C}$ for which these structure theorems hold; see Question 9.4 for a specific question on how this might be the case.
@@ 1.1 Braid groups and the Artin representation
[…]
@@
[…]
@@ 2.2 The tensor algebra and its dual
$[a_{1}|\dots|a_{p}]\star[b_{1}|\dots|b_{q}]=\overline{\mathbf{S}}_{p,q}\cdot[a_{1}|\dots|a_{p}|b_{1}|\dots|b_{q}].$
[…]
@@ 3.1 Definition and examples
[…]
2=123121234123121234
Figure 1: Cabling a 2-strand braid onto the second strand (as read from the top) in a 3-strand braid.
@@
    (Associativity) For each $m$, $n$, and $p$, and indices $1\leq j\leq m$ and $11\leq i\leq m+n-$ 1, there is an equality of the map
@@
    • $\circ_{j+p-1}(\circ_{i}\otimes\mbox{id}_{\mathcal{C}(n)})(\mbox{id}_{\mathcal{C}(m)}\otimes\tau)$, if $11\leq i\leq j-$ 1.
@@
[…]
@@
    $[-,-]_{\sigma}:=\mu-\mu\circ\sigma.$
@@
[…]
@@ 3.2 Algebras for braided operads
The data of an algebra over $\operatorname{BrAss}$ in $\mathbf{C}$ is the same as that of a unital, associative algebra object $A\in\mathbf{C}$.
@@
Similarly, an algebra over $\operatorname{BrCom}$ is a braided commutative, unital, associative algebra. That is, the multiplication $\mu$ on $A$ satisfies $\mu=\mu\circ\sigma$.
These examples highlight the fact that algebras for unitary operads are equipped with elements that serve as units for the operations that the operad defines. In contrast, algebras for the reduced forms $\mathcal{C}_{>0}$ of these operads should be thought of as not-necessarily-unital $\mathcal{C}$-algebras.
@@
    The free braided Lie algebra $\operatorname{BrLie}_{*}[V]$ is the smallest subspace of $T(V)$ which contains $V$ and is closed under $[-,-]_{\sigma}$ and the $B_{n}$ action in degree $n$.
@@ 3.3 Modules for algebras over an operad
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Ideals in $\operatorname{BrAss}$ and $\operatorname{BrCom}$-algebras are the same thing as two-sided ideals in the usual sense. An ideal $I$ in a $\operatorname{BrLie}$-algebra $L$ is a sub-object which is closed under $[\ell,-]$ for every $\ell\in L$.
@@ 3.5 Enveloping algebras
    where $[A,A]_{\sigma}$ is the two-sided ideal generated by the braided brackets $[a,b]_{\sigma}$.
@@
In a similar fashion, one may more generally show that $U_{\mathcal{C}}^{\mathcal{D}}(L/I)\cong U_{\mathcal{C}}^{\mathcal{D}}(L)/(I)$ for any $\mathcal{C}$-ideal $I\leq L$.
@@ 4.1 Pro-operads and profinite completion
[…]
@@
${\operatorname{BrOp}(\operatorname{pro}\!-\mathbf{D})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widehat{}}$${\widehat{\operatorname{BrOp}}(\operatorname{pro}\!-\mathbf{D})}$
@@
[…]
@@
[…]
@@ 4.2 Algebras for the profinite completion of an operad
Now consider an operad $\mathcal{O}$ in $\mathbf{D}$ and its profinite completion $\widehat{\mathcal{O}}$ in $\operatorname{pro}\!-\mathbf{D}$. The unit of the adjunction in Proposition 4.4 ensures that an $\widehat{\mathcal{O}}$-algebra $A$ may be regarded as a $\mathcal{O}$-algebra, but the converse need not be the case. Such an algebra structure would amount to a family of liftings
[…]
@@ 4.3 The value of the Artin functor for subsequences of $\widehat{\operatorname{BrAss}}$
[…]
Since the characteristic of $k$ is 00 and $Q$ is finite, if $M$ is a $k[Q]$-module, the operation $-\otimes_{k[Q]}M$ is exact, yielding an exact sequence
[…]
@@ 5.1 The monad of primitives
$PT$ is a monad, and if $A$ is a braided Hopf algebra, $P(A)$ is an algebra for $PT$.
@@
[…]
@@ 5.2 The braided primitive operad
[…]
@@
We conclude by observing that this isomorphism equates the monadic compositions of $PT$ and $\operatorname{BrPrim}_{*}[-]$. To see this, note that both are subfunctors of the tensor algebra functor. The previous argument shows that these are in fact the same subfunctor of $T(-)$. In both cases, the monadic compositions are defined as a (co)restriction of the tensor algebra monad, and hence agree.
@@ 5.3 The structure of $\operatorname{BrPrim}$
Further, if $m|n$, the image of $b_{2n}$ in $k[\mathbb{Z}/2m]$ is precisely $b_{2m}$. Thus the collection assembles into a well-defined element $b$ in the inverse system, $\operatorname{BrPrim}(2)$. Since each map in the system is an isomorphism, the system is pro-constant, rank 1, and clearly generated by $b$. By definition, the action of $B_{2}$ on $\operatorname{BrPrim}(2)$ equalizes $\sigma$ and $-1$; this is the sign representation.
@@
[…]
@@
[…]
@@
[…]

### v2

@@ Abstract
The cross terms cancel, showing that [a,b]=abba[a,b]=ab-ba is primitive.
@@
Consequently, the same argument that $P(A)$ is a Lie algebra fails in the braided context. A great deal of the technical work in this paper is spent developing the notion of a braided primitive algebra [^1] The term “braided Lie algebra” is already in use in multiple (inequivalent) forms; see section 10 for a review of some of these terms and their relation to the constructions in this paper., which encapsulates the algebraic properties of $P(A)$. This structure is encoded by the data of an action of a braided operad which we notate as $\operatorname{BrPrim}$ and define in Section 5.2.
@@
The primitives $P(A)$ sit inside a larger space $P_{\mathfrak{W}_{*}}(A)\subseteq A$ whose definition (in section 7.1) requires some care, but can loosely be described as “all of the elements of $A$ that become primitive after iteratively quotienting by decomposable primitives in AA” (see Proposition 9.2 for a more precise description). It supports an algebraic structure we are calling [^2] This term is also used in the setting of $C^{*}$-algebras; these are different notions, and it should be clear from context which is meant. a Woronowicz algebra. The corresponding operad is denoted $\mathfrak{W}_{*}$; it is defined in section 5.5 in terms of the “quantum symmetrizer” which gives the relations in Nichols algebras. It is larger and even less well-understood than $\operatorname{BrPrim}$.
@@
If $L$ is a Woronowicz algebra equipped with a diagonal, there is a filtration on $U_{\mathfrak{W}_{*}}(L)$ by powers of the primitives $P(L)=P(U_{\mathfrak{W}_{*}}(L))$. Unlike the classical setting, it is possible that more primitives are produced in the associated graded Hopf algebra $U_{\mathfrak{W}_{*}}(L)^{\operatorname{gr}}$. We may, however, iterate this procedure, producing a graded Hopf algebra $U_{\mathfrak{W}_{*}}(L)^{(\infty)}$. Our analogue of the Poincaré-Birkhoff-Witt theorem is:
@@
We actually prove versions of Theorems 1 and 2 for a large class of operads $\mathcal{C}$; see Theorem 7.10 and Proposition 8.14 for precise statements. The Woronowicz ideal $\mathfrak{W}$ is the smallest operad $\mathcal{C}$ that we have identified for which these results hold, but still very large. It is very plausible that there is a class of simpler (or at least smaller) operads $\mathcal{C}$ for which these structure theorems hold; see Question 9.4 for specifics on how this might be the case.
@@ 1.1 Braid groups and the Artin representation
Recall that the $n\textsuperscript{th}$ braid group $B_{n}$ is presented by
@@
Pictorially, $\sigma_{i}$ is the braid which permutes the $i\textsuperscript{th}$ and $i+1\textsuperscript{st}$ endpoints, where the strand beginning at the latter passes over the one beginning at the former. Additionally, we interpret $B_{0}$ and $B_{1}$ to be the trivial group.
@@ 2.2 The tensor algebra and its dual
$a_{1}|\dots|a_{p}]\star[b_{1}|\dots|b_{q}]=\overline{\mathbf{S}}_{p,q}\cdot[a_{1}|\dots|a_{p}|b_{1}|\dots|b_{q}].$
If $\mathbf{C}$ is rigid, and $V$ an object of $\mathbf{C}$, the $n\textsuperscript{th}$ tensor power $(V^{*})^{\otimes n}\cong(V^{\otimes n})^{*}$ of its dual $V^{*}$ carries two actions of $B_{n}$: a left action naturally arising from the structure of the category, and a right action as the adjoint of the left action on $V^{\otimes n}$. It is straightforward to see that these actions are thrown onto each other by inversion in $B_{n}$. As such, one may identify $T^{\operatorname{co}}(V)$ as the (graded) dual braided Hopf algebra to $T(V^{*})$:
@@ 3.1 Definition and examples
For $\gamma\in B_{m}$ and $\rho\in B_{n}$, $\gamma\circ_{i}\rho$ is obtained by fattening the $i\textsuperscript{th}$ strand of $\gamma$ to a tubular neighborhood, and filling that tube with $\rho$. This is not a group homomorphism, but becomes one when restricted $B_{m}^{i}\times B_{n}$, where $B_{m}^{i}\leq B_{m}$ is the subgroup fixing $i\in\{1,\dots,m\}$ under the homomorphism $B_{m}\to S_{m}$.
2=$123121234$
Figure 1: Cabling a 2-strand braid onto the second strand (with strands read from the top) in a 3-strand braid.
@@
    (Associativity) For each $m$, $n$, and $p$, and indices $1\leq j\leq m$ and $1\leq i\leq m+n-1$, there is an equality of the map
@@
    • $\circ_{j+p-1}(\circ_{i}\otimes\mbox{id}_{\mathcal{C}(n)})(\mbox{id}_{\mathcal{C}(m)}\otimes\tau)$, if $1\leq i\leq j-1$.
@@
One is meant to regard $\mathcal{C}(n)$ as an abstract family of $n$-ary operations, and the operation $\circ_{i}:\mathcal{C}(m)\otimes\mathcal{C}(n)\to\mathcal{C}(m+n-1)$ as inserting the second $n$-ary factor into the $i\textsuperscript{th}$ input of the first $m$-ary factor. The above is a braided adaptation of Markl’s definition [Mar96] of a “pseudo-operad,” rather than May’s original definition [May72] (which would give precisely Fiedorowicz’s definition of a braided operad). These are equivalent if the operad is unital, and Markl’s definition is slightly more general if not. In particular, the composite of all of the $\circ_{i}$ defines a family of maps
@@
    $_{\sigma}:=\mu-\mu\circ\sigma.$
@@
    The $n\textsuperscript{th}$ braid group acts on $\operatorname{End}_{X}(n)$ by precomposition, and the structure map $\circ_{i}$ comes from substitution of $g$ into the $i\textsuperscript{th}$ tensor factor:
@@ 3.2 Algebras for braided operads
The data of an algebra over $\operatorname{BrAss}$ in $\mathbf{C}$ is the same as that of an associative algebra object $A\in\mathbf{C}$ with unit.
@@
Similarly, an algebra over $\operatorname{BrCom}$ is a braided commutative, associative algebra with unit. That is, the multiplication $\mu$ on $A$ satisfies $\mu=\mu\circ\sigma$.
These examples highlight the fact that algebras for unitary operads are equipped with elements that serve as units for the operations that the operad defines. In contrast, algebras for the reduced forms $\mathcal{C}_{>0}$ of these operads should be thought of as $\mathcal{C}$-algebras which are not required to have such a unit.
@@
    The free algebra $\operatorname{BrLie}_{*}[V]$ for the unital form of the braided Lie operad is the smallest subspace of $T(V)$ which contains $V$ and is closed under $_{\sigma}$ and the $B_{n}$ action in degree $n$.
@@ 3.3 Modules for algebras over an operad
If $\mathcal{C}$ is an operad and $A$ is a $\mathcal{C}$-algebra in $\mathbf{C}$, an $A$-module $M$ is an object of $\mathbf{C}$ equipped with maps [^11] Recall in this formula that $B_{n}^{i}<B_{n}$ is the subgroup where the $i\textsuperscript{th}$ strand is pure.
@@
${(\mathcal{C}(m)\otimes\mathcal{C}(n))\otimes A^{\otimes n+m-2}\otimes M\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\circ_{i}\otimes 1}$$\scriptstyle{1\otimes\tau\otimes 1}$${\mathcal{C}(m+n-1)\otimes A^{\otimes n+m-2}\otimes M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda}$${\mathcal{C}(m)\otimes A^{\otimes{i-1}}\otimes(\mathcal{C}(n)\otimes A^{\otimes n})\otimes A^{\otimes m-i-1}\otimes M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{1\otimes\theta\otimes 1}$${\mathcal{C}(m)\otimes A^{\otimes m-1}\otimes M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda}$${M}$
@@
${(\mathcal{C}(m)\otimes\mathcal{C}(n))\otimes A^{\otimes n+m-2}\otimes M\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\circ_{m}\otimes 1}$$\scriptstyle{1\otimes\tau\otimes 1}$${\mathcal{C}(m+n-1)\otimes A^{\otimes n+m-2}\otimes M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda}$${\mathcal{C}(m)\otimes A^{\otimes{m-1}}\otimes(\mathcal{C}(n)\otimes A^{\otimes n-1}\otimes M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{1\otimes\lambda}$${\mathcal{C}(m)\otimes A^{\otimes m-1}\otimes M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda}$${M.}$
@@
${\mathbf{1}\otimes M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\mathcal{C}(1)\otimes M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda}$${M}$
@@
Ideals in $\operatorname{BrAss}$ and $\operatorname{BrCom}$-algebras are the same thing as two-sided ideals in the usual sense. An ideal $I$ in a $\operatorname{BrLie}$-algebra $L$ is a sub-object which is closed under $\ell$,-] for every $\ell\in L$.
@@ 3.5 Enveloping algebras
    where $A,A]_{\sigma}$ is the two-sided ideal generated by the braided brackets $a,b]_{\sigma}$.
@@
In a similar fashion, one may more generally show that $U_{\mathcal{C}}^{\mathcal{D}}(L/I)\cong[U_{\mathcal{C}}^{\mathcal{D}}(L)]/(I)$ for any $\mathcal{C}$-ideal $I\leq L$.
@@ 4.1 Pro-operads and profinite completion
which carries an operad $\mathcal{C}$ to the pro-operad whose $n\textsuperscript{th}$ term is the constant inverse system on $\mathcal{C}(n)$. This is induced by the symmetric monoidal functor $\operatorname{const}:\mathbf{D}\to\operatorname{pro}\!-\mathbf{D}$ which carries objects to constant inverse systems. Since it is monoidal, it indeed carries operads to operads.
@@
${\operatorname{BrOp}(\operatorname{pro}\!-\mathbf{D})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}^\scriptstyle{\widehat{}}$${\widehat{\operatorname{BrOp}}(\operatorname{pro}\!-\mathbf{D})}$
@@
are defined by descending the structure maps (4.7) on $\mathcal{C}$ through the coinvariants. Note that the maps in (4.7) are equivariant with respect to the cabling homomorphism on the braid groups $\circ_{i}:B_{n}^{i}\times B_{m}\to B_{n+m-1}$ (recall that in $B_{n}^{i}$, the $i\textsuperscript{th}$ strand is pure). Thus the maps in (4.8) are well defined if
@@
The profinite braided associative operad $\widehat{\operatorname{BrAss}}$ is the profinite completion of $\operatorname{BrAss}$. Its $n\textsuperscript{th}$ term is the inverse system $\widehat{\operatorname{BrAss}}(n)=\{k[Q],\;Q\in\operatorname{Quot}_{n}\}$ of group rings of finite quotients $Q$ of $B_{n}$.
@@ 4.2 Algebras for the profinite completion of an operad
Now consider an operad $\mathcal{O}$ in $\mathbf{D}$ and its profinite completion $\widehat{\mathcal{O}}$ in $\operatorname{pro}\!-\mathbf{D}$. The unit of the adjunction in Proposition 4.4 ensures that an $\widehat{\mathcal{O}}-$algebra AA may be regarded as a $\mathcal{O}$-algebra, but the converse need not be the case. Such an algebra structure would amount to a family of liftings
${\operatorname{Hom}_{\mathbf{C}}(A^{\otimes n},A)^{\ker_{Q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\subseteq}$${\mathcal{O}(n)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\theta}$${\operatorname{Hom}_{\mathbf{C}}(A^{\otimes n},A)}$
@@ 4.3 The value of the Artin functor for subsequences of $\widehat{\operatorname{BrAss}}$
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\ker(f)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${k[Q]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${k[Q]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{cok}(f)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
Since the characteristic of $k$ is 0 and $Q$ is finite, if $M$ is a $k[Q]$-module, the operation $\otimes_{k[Q]}M$ is exact, yielding an exact sequence
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\ker(f)\otimes_{k[Q]}M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{cok}(f)\otimes_{k[Q]}M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0,}$
@@ 5.1 The monad of primitives
PTPT is a monad, and if $A$ is a braided Hopf algebra, $P(A)$ is an algebra for $PT$.
@@
${T(P(T(P(A))))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{T(P(M_{A}))}$$\scriptstyle{M_{T(P(A))}}$${T(P(A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{M_{A}}$${T(P(A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{M_{A}}$${A.}$
@@ 5.2 The braided primitive operad
Let $\alpha\in\operatorname{BrPrim}_{*}(n)$ and $\beta\in\operatorname{BrPrim}_{*}(m)$; we must show that the $i\textsuperscript{th}$ cabling $\alpha\circ_{i}\beta$ is in $\operatorname{BrPrim}_{*}(n+m-1)$. By the previous theorem, it is enough to show that for any finitely braided Hopf algebra $A$ and elements $a_{1},\dots,a_{i-1},b_{1},\dots,b_{m},a_{i+1},\dots a_{n}\in P(A)$,
@@
We conclude by observing that this isomorphism equates the monadic compositions of $PT$ and $\operatorname{BrPrim}_{*}[-]$. To see this, note that both are subfunctors of the tensor algebra functor. The previous argument shows that these are in fact the same subfunctor of T()T(-). In both cases, the monadic compositions are defined as a (co)restriction of the tensor algebra monad, and hence agree.
@@ 5.3 The structure of $\operatorname{BrPrim}$
Further, if $m|n$, the image of $b_{2n}$ in $k[\mathbb{Z}/2m]$ is precisely $b_{2m}$. Thus the collection assembles into a well-defined element $b$ in the inverse system, $\operatorname{BrPrim}(2)$. Since each map in the system is an isomorphism, the system is pro-constant, rank 1, and clearly generated by $b$. By definition, the action of $B_{2}$ on $\operatorname{BrPrim}(2)$ equalizes $\sigma$ and 1-1; this is the sign representation.
@@
Consider a braided operad $\mathcal{C}$ and a left module $\mathcal{M}$ for $\mathcal{C}$. Assume that $\mathcal{C}(0)=0=\mathcal{C}(1)$ and $\mathcal{M}(0)=0$. The indecomposables of $\mathcal{M}$ are given by the braided sequence $Q_{\mathcal{C}}(\mathcal{M})$ whose $n\textsuperscript{th}$ term is the cokernel of the map
@@
Since $\operatorname{char}(k)=0$ we may basechange to a field containing a primitive $n\textsuperscript{th}$ root of unity, $\zeta$. Then the braided vector space $V=k$ with braiding given by multiplication with $\zeta$,
@@
To see this, we note that Ree’s Theorem 2.2 [Ree58] implies that the $n\textsuperscript{th}$ term $\operatorname{Lie}(n)$ of the Lie operad is precisely